We were able to resolve the AOM double-pass issue by enlarging the beam going into it. We had originally tried using the smallest practical waist (~ 70 um) because the jitter suppression scheme only works perfectly--i.e. beam ROC = mirror ROC with the mirror exactly R away from the AOM--in the limit where zR --> 0, or w0 = 0. It appears that this caused the beam to diverge too quickly within the AOM, which is pretty long. We have changed the waist size to ~200 um at the expense of moving the mirror in to z = 15.5 cm from the AOM (as opposed to z = R = 30 cm as it was before). We are now able to isolate a decent-looking double-pass beam, which we verified by measuring its power. Hopefully the loss in jitter suppression will be insignificant as we only expect the AOM to shift by hundreds of Hz or so, corresponding to a very small angular displacement.
After this progress this morning, I tried to realign the cavity and get the original unmodulated direction to lock, but I am having quite some difficulty. I can align it and isolate the 00 mode pretty well, and there appears to be a decent amount of power emerging at the transmitted end on resonance (as seen on IR card), but the thing just doesn't want to lock. I have tried to adjust the RF phase to no avail. I suspect it has something to do with a polarization issue with the BS I am using to screen some power from the REFL beam before the PD, but I can't figure it out. I will try again tomorrow morning.
Once this direction locks again, we can feed in the modulated beam around the other direction, and we should have all the necessary equipment to set up the second loop. I spoke with Frank today, and I believe we will need a Marconi to set up a PLL at the output to do the frequency readout. We will figure this out when it comes up, I guess.